# Xingwang Li, , and Arumugam Nallanathan

### v1

@@ III-B2 Linear Estimation Module
$\mathbf{H}_{\text{linear}}=(\mathbf{Y}_{\text{rec}}-\mathbf{B})(\mathbf{S}^{T})^{\dagger}.$ (19)
@@ III-B3 Residual Correction Module
To remove the contained residual errors in $\mathbf{H}_{\text{linear}}$, we develop a Transformer-based residual correction module, termed $\text{Former}_{\phi}$ with trainable weights $\phi$}, which can be expressed as
$\hat{\mathbf{H}}^{(t)}=\mathbf{H}_{\text{linear}}+\text{Former}_{\phi}(\mathbf{H}_{\text{linear}}),$ (20)
@@
To capture the complex non-local dependencies within the MIMO channel matrix which arise from physical phenomena like clustered multipath propagation, we implement $\text{Former}_{\phi}$ using a channel Transformer based on the self-attention mechanism [11]. Let the input to the module be the linear channel estimate $\mathbf{H}_{\text{linear}}\in\mathbb{C}^{M\times K}$. The function of the residual correction module, $\text{Former}_{\phi}(\mathbf{H}_{\text{linear}})$, is defined by the following sequence of operations.
@@
  • Tokenization: First, the complex input matrix is decomposed into its real and imaginary components, denoted by operators $\Re(\cdot)$ and $\Im(\cdot)$, respectively, and reshaped into a sequence of $K$ tokens, i.e., $\mathbf{X}_{\text{in}}=\text{stack}(\Re(\mathbf{H}_{\text{linear}}),\Im(\mathbf{H}_{\text{linear}}))\in\mathbb{R}^{2\times M\times K}$ is the initial tensor representation. This tensor is then reshaped into a token sequence $\mathbf{T}=\text{Reshape}(\mathbf{X}_{\text{in}})\in\mathbb{R}^{K\times 2M}$.
@@
  • Embedding: The token sequence $\mathbf{T}$ is projected into a higher-dimensional latent space, and positional information is added. This forms the input to the first Transformer layer $\mathbf{Z}^{(0)}\in\mathbb{R}^{K\times d_{\text{model}}}$, which can be expressed as
    $\mathbf{Z}^{(0)}=\mathbf{T}\mathbf{W}_{\text{proj}}+\mathbf{P}_{\text{pos}},$ (21)
@@
    where $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{2M\times d_{\text{model}}}$ denotes the learnable weight matrix of the input projection layer, and $\mathbf{P}_{\text{pos}}\in\mathbb{R}^{K\times d_{\text{model}}}$ is the learnable positional embedding matrix.
@@
  • Transformer Encoder Blocks: The embedded sequence $\mathbf{Z}^{(0)}$ is then processed by a stack of $L$ identical Transformer encoder layers. For each layer $l=1,\dots,L$, the multi-head self-attention (MHSA) is carried out, i.e., $\mathbf{A}^{(l)}=\text{MHSA}(\text{LN}(\mathbf{Z}^{(l-1)}))+\mathbf{Z}^{(l-1)}$. Then, a feed-forward network (FFN) is utilized to obtain the feature tensor $\mathbf{Z}^{(l)}=\text{FFN}(\text{LN}(\mathbf{A}^{(l)}))+\mathbf{A}^{(l)}$, where $\text{LN}(\cdot)$ denotes layer normalization. The output of the final layer is $\mathbf{Z}^{(L)}\in\mathbb{R}^{K\times d_{\text{model}}}$.
@@
  • De-Tokenization: The processed sequence $\mathbf{Z}^{(L)}$ is projected back to the original token dimension and reassembled into the matrix format. The output token sequence is given by $\mathbf{T}_{\text{out}}=\text{FFN}_{\text{out}}(\mathbf{Z}^{(L)})\in\mathbb{R}^{K\times 2M}$, which is reshaped back into a tensor $\mathbf{X}_{\text{res}}\in\mathbb{R}^{2\times M\times K}$. Finally, the two channels of $\mathbf{X}_{\text{res}}$ are combined to form the complex residual matrix $\mathbf{H}_{\text{res}}=\text{Former}_{\phi}(\mathbf{H}_{\text{linear}})$. Let $\mathbf{X}_{\text{res}}[0]$ and $\mathbf{X}_{\text{res}}[1]$ be the first and second channels of $\mathbf{X}_{\text{res}}$, respectively, we have $\text{Former}_{\phi}(\mathbf{H}_{\text{linear}})=\mathbf{X}_{\text{res}}[0]+i\cdot\mathbf{X}_{\text{res}}[1]$.
@@
The entire $T_{\text{UR}}$-layer URformer network is trained end-to-end with the normalized mean squared error (NMSE) loss function, which is given by
@@
IV Simulation Results
Table I summarizes the main system parameters of the considered RAQR and training hyper-parameter setups of the proposed URformer. We compare the proposed URformer against several baseline methods: the classic GS and EM-GS algorithms, as well as two black-box deep learning approaches, i.e., the convolutional neural neiwork (CNN) and the standard Transformer architecture. For fair comparison, the number of iterations for the classic algorithms is set to be $T_{\text{GS}}=100$ to provide superior estimation accuracy for iterative algorithms.
TABLE I: Simulation Parameters
Parameter Value
@@ IV Simulation Results
Number of Transformer encoders $L$ 3
@@
Fig. 3 illustrates the NMSE performance of all methods as a function of the received SNR that is defined as $\mathrm{SNR}={\mathbb{E}\left(|\mathbf{H}\mathbf{s}_{p}|^{2}\right)}/{\mathbb{E}\left(|\mathbf{w}_{p}|^{2}\right)}$, where the reference-to-signal ratio (RSR) is set to $\mathrm{RSR}={\mathbb{E}\left(|\mathbf{b}|^{2}\right)}/{\mathbb{E}\left(|\mathbf{H}\mathbf{s}_{p}|^{2}\right)}=10$ dB. The results clearly demonstrate the superiority of the proposed URformer. In the challenging low-SNR regime, e.g., 0 dB, where the classic GS and EM-GS algorithms struggle significantly, URformer achieves a remarkable performance gain. This highlights the robustness of its learned modules, particularly the neural filter and Transformer block, which can effectively extract channel features even from very noisy measurements. In contrast, the black-box CNN and Transformer models fail completely, yielding a high error floor across all SNRs. This indicates their inability to solve the complex biased phase retrieval problem without the structural guidance of the underlying physical model.
@@
Fig. 4 shows the NMSE performance versus the number of pilot transmissions $P$, evaluated at a fixed SNR of 5 dB. This analysis is crucial for assessing the pilot efficiency of each algorithm. As expected, the performance of all methods improves with an increased number of pilots. However, the proposed URformer exhibits a much steeper performance curve, indicating a significantly higher pilot efficiency. For instance, URformer achieves an NMSE of -25 dB with only $P=15$ pilots. To reach a similar performance level, the classic algorithms would require a far greater number of pilots. This superior pilot efficiency allows for a reduction in pilot overhead, thereby increasing the spectral efficiency of RAQR-based communication systems.
@@ V Conclusion
In this work, we addressed the challenging problem of channel estimation for RAQRs, which are characterized by a unique, non-linear biased phase retrieval signal model. Moving beyond the limitations of classic iterative algorithms and data-agnostic black-box models, we proposed a novel model-driven URformer architecture by unrolling a stabilized version of the physically-grounded EM-GS algorithm. The each unrolling layer of the proposed URformer is composed of a learnable neural filter, a stabilizing gate, and a channel Transformer module, allowing it to overcome the limitations of classic iterative methods. Numerical results demonstrated that the proposed URformer significantly outperformed the existing algorithms and showed substantially higher pilot efficiency. Furthermore, its vast performance advantage over black-box CNN and Transformer baselines validates the power of the model-driven deep learning paradigm.

### v2

@@ III-B2 Linear Estimation Module
$\mathbf{H}_{\text{linear}}^{(t-1)}=(\mathbf{Y}_{\text{rec}}^{(t-1)}-\mathbf{B})(\mathbf{S}^{T})^{\dagger}.$ (17)
@@ III-B3 Residual Correction Module
To remove the contained residual errors in $\mathbf{H}_{\text{linear}}^{(t-1)}$, we develop a Transformer-based residual correction module, termed $\text{Former}_{{\phi}^{(t-1)}}$ with trainable weights ${\phi}^{(t-1)}$, which can be expressed as
$\hat{\mathbf{H}}^{(t)}=\mathbf{H}_{\text{linear}}^{(t-1)}+\text{Former}_{{\phi}^{(t-1)}}(\mathbf{H}_{\text{linear}}^{(t-1)}),$ (18)
@@
To capture the complex non-local dependencies within the MIMO channels which arise from clustered multipath propagation, we develop a channel Transformer module, i.e., $\text{Former}_{{\phi}^{(t-1)}}$, based on the self-attention mechanism [11, 12]. Let the input to the module $\text{Former}_{{\phi}^{(t-1)}}$ be the linear channel estimate $\mathbf{H}_{\text{linear}}^{(t-1)}\in\mathbb{C}^{M\times K}$. The residual correction module is defined by the following sequence of operations.
@@
  • Tokenization: First, the complex input matrix is decomposed into its real and imaginary components, denoted by operators $\Re(\cdot)$ and $\Im(\cdot)$, respectively, and reshaped into a sequence of $K$ tokens, i.e., $\mathbf{X}_{\text{in}}^{(t-1)}=\{\Re(\mathbf{H}_{\text{linear}}^{(t-1)}),\Im(\mathbf{H}_{\text{linear}}^{(t-1)})\}\in\mathbb{R}^{2\times M\times K}$ is the initial tensor representation. This tensor is then reshaped into a token sequence $\mathbf{T}_{\text{in}}^{(t-1)}=\text{Reshape}(\mathbf{X}_{\text{in}}^{(t-1)})\in\mathbb{R}^{K\times 2M}$.
@@
  • Embedding: The token sequence $\mathbf{T}_{\text{in}}^{(t-1)}$ is projected into a higher-dimensional latent space, and positional information is added. This forms the input to the first Transformer layer $\mathbf{V}^{(0)}\in\mathbb{R}^{K\times d_{\text{model}}}$, which can be expressed as
    $\mathbf{V}^{(0)}=\mathbf{T}_{\text{in}}^{(t-1)}\mathbf{W}+\mathbf{P},$ (19)
@@
    where $\mathbf{W}\in\mathbb{R}^{2M\times d_{\text{model}}}$ denotes the learnable weight matrix of the input projection layer, and $\mathbf{P}\in\mathbb{R}^{K\times d_{\text{model}}}$ is the learnable positional embedding matrix.
    Figure 2: Proposed URformer architecture for channel estimation.
@@
  • Transformer Encoder Blocks: The embedded sequence $\mathbf{V}^{(0)}$ is then processed by a stack of $L_{\text{enc}}$ identical Transformer encoder layers. For each layer $l=1,\dots,L_{\text{enc}}$, the multi-head self-attention (MHSA) is carried out, i.e., $\mathbf{A}^{(l)}=\text{MHSA}(\text{LN}(\mathbf{V}^{(l-1)}))+\mathbf{V}^{(l-1)}$. Then, a feed-forward network (FFN) is utilized to obtain the feature tensor $\mathbf{V}^{(l)}=\text{FFN}(\mathbf{A}^{(l)})+\mathbf{A}^{(l)}$. The output of the final layer is $\mathbf{V}^{(L_{\text{enc}})}\in\mathbb{R}^{K\times d_{\text{model}}}$.
@@
  • De-Tokenization: The processed sequence $\mathbf{V}^{(L_{\text{enc}})}$ is projected back to the original token dimension. The output token sequence is given by $\mathbf{T}_{\text{out}}^{(t-1)}=\text{FFN}_{\text{out}}(\mathbf{V}^{(L_{\text{enc}})})\in\mathbb{R}^{K\times 2M}$, which is then reshaped into a tensor $\mathbf{X}_{\text{res}}^{(t-1)}\in\mathbb{R}^{2\times M\times K}$. Finally, the two channels of $\mathbf{X}_{\text{res}}^{(t-1)}$ are combined to form the complex residual matrix $\mathbf{H}_{\text{res}}^{(t-1)}=\text{Former}_{{\phi}^{(t-1)}}(\mathbf{H}_{\text{linear}}^{(t-1)})$.
@@
The $T_{\text{UR}}$-layer URformer is trained with the normalized mean squared error (NMSE) loss function, which is given by
@@
The classic GS and EM-GS algorithms have a complexity of $\mathcal{O}(T_{\mathrm{GS}}MKP)$, while the proposed URformer has a complexity of $\mathcal{O}(T_{\mathrm{UR}}(MKP+L_{\mathrm{enc}}Kd_{\mathrm{model}}(K+d_{\mathrm{model}})))$. Although the URformer incurs an additional $L_{\mathrm{enc}}Kd_{\mathrm{model}}(K+d_{\mathrm{model}})$ term per layer due to the self-attention mechanism, it achieves convergence with significantly fewer unrolled layers ($T_{\mathrm{UR}}\ll T_{\mathrm{GS}}$), thereby reducing the overall computational burden.
IV Numerical Results
In this work, we consider a Cesium (Cs) vapor cell-based RAQR and the four-level ladder scheme is set to $6S_{1/2}\to 6P_{3/2}\to 47S_{1/2}\to 47P_{1/2}$ for a resonant millimeter-wave frequency $f_{c}=37.406$ GHz [13]. The specific parameter configurations of RAQR are obtained with the Alkali Rydberg Calculator (ARC) package [14]. The proposed URformer is trained by the Adam optimizer with a cosine annealing learning rate scheduler over $T_{\text{max}}=50$ epochs. Table I summarizes the main system parameters of the considered RAQR and training hyper-parameter setups of the proposed URformer. We compare the proposed URformer against several baseline methods: the classic GS and EM-GS algorithms, as well as two black-box deep learning approaches, i.e., the convolutional neural network (CNN) and the standard Transformer architecture [12]. For fair comparison, the number of iterations for the classic algorithms is set to be $T_{\text{GS}}=100$ to provide superior estimation accuracy for iterative algorithms.
TABLE I: Parameter Setups for Channel Estimation in RAQRs
Parameters Values
@@ IV Numerical Results
Number of Transformer encoders $L_{\text{enc}}$ 3
@@
Number of training batchsize 32
Initial learning rate $1\times 10^{-3}$
@@
Fig. 3 illustrates the NMSE performance of all methods as a function of the received SNR that is defined as $\mathrm{SNR}={\mathbb{E}\left(|\mathbf{H}\mathbf{s}_{p}|^{2}\right)}/{\mathbb{E}\left(|\mathbf{w}_{p}|^{2}\right)}$, where the reference-to-signal ratio (RSR) is set to $\mathrm{RSR}={\mathbb{E}\left(|\mathbf{b}|^{2}\right)}/{\mathbb{E}\left(|\mathbf{H}\mathbf{s}_{p}|^{2}\right)}=10$ dB and the number of pilots is set to $P=20$. The results clearly demonstrate the superiority of the proposed URformer. In the challenging low-SNR regime, e.g., 0 dB, where the classic GS and EM-GS algorithms struggle significantly, URformer achieves a remarkable performance gain. This highlights the robustness of its learned modules, particularly the neural filter and Transformer block, which can effectively extract channel features even from very noisy measurements. In contrast, the black-box CNN and Transformer models fail completely, yielding a high error floor across all SNRs. This indicates their inability to solve the complex phase retrieval problem without the structural guidance of the underlying physical model.
@@
Fig. 4 shows the NMSE performance versus the number of pilot transmissions $P$, evaluated at a fixed SNR of 5 dB. This analysis is crucial for assessing the pilot efficiency of each algorithm. As expected, the performance of all methods improves with an increased number of pilots. However, the proposed URformer exhibits a much steeper performance curve, indicating a significantly higher pilot efficiency. For instance, URformer achieves an NMSE of -20 dB with only $P=15$ pilots. This superior pilot efficiency allows for a reduction in pilot overhead, thereby increasing the spectral efficiency of RAQR-based communication systems.
@@ V Conclusion
In this letter, we investigated the channel estimation for RAQRs, which are characterized by a non-linear biased phase retrieval signal model from the measurement of holographic snapshots. Moving beyond the limitations of classic iterative algorithms and data-agnostic black-box models, we proposed a novel model-driven URformer architecture by unrolling a stabilized version of the classic EM-GS algorithm. Each unrolling layer of the proposed URformer is composed of a learnable neural filter, a stabilizing gate, and a channel Transformer module, allowing it to overcome the limitations of classic iterative methods. Numerical results demonstrated that the proposed URformer significantly outperformed the existing algorithms and showed substantially higher pilot efficiency.